OpenStudy (anonymous):

A Norman window has the shape of a semicircle atop a rectangle so that the diameter of the semicircle is equal to the width of the rectangle. What is the area of the largest possible Norman window with a perimeter of 31 feet?

OpenStudy (phi):

looks familiar

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

hey mubzz can you help me pls

OpenStudy (anonymous):

Area of rectangle = (2r)*h = 2rh Area of semicircle = 1/2 pi r^2 So we want to maximize A = 2rh + 1/2 pi r^2 but there are two variables, so we'll first try to eliminate one of them. Note that perimeter = 2h + 2r + pi r = 31 so h = 31 - r - pi/2 r = 31 - (1 + pi/2)r Substitute that into our equation for area to get: A = 2r(31-(1+pi/2)r) + 1/2 pi r^2 = 62r - 2r^2 - 1/2 pi r^2 Take the derivative: dA/dr = 62 - 4r - pi r and set it equal to zero: 0 = 62 - 4r - pi r (4 + pi) r = 62 r = 62 / (4 + pi) Now that you have the radius, you can find the area A by using the equation A = 62r - 2r^2 - 1/2 pi r^2

OpenStudy (anonymous):

my program wants a numerical value

OpenStudy (anonymous):

sooo

OpenStudy (anonymous):

i just substitute r?

OpenStudy (anonymous):

just set r = 31 feet ans solve it

OpenStudy (anonymous):

A = 62r - 2r^2 - 1/2 pi r^2 = 62*31 - 2(31)^2 - (1/2)*pi*(31)^2 = 1922 - 1922 - 1510.14 |A| = 1510.14

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

my program says no

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

this person probably know how to solve this problem

OpenStudy (anonymous):

I'm sorry but I really can't get it

OpenStudy (anonymous):

ok what about this A box is to be made out of a 8 by 16 piece of cardboard. Squares of equal size will be cut out of each corner, and then the ends and sides will be folded up to form a box with an open top. Find the length L, width W, and height H of the resulting box that maximizes the volume. (Assume that W≤L).

OpenStudy (phi):

@smr_kml Note that perimeter = 2h + 2r + pi r = 31 so h = 31 - r - pi/2 r = 31 - (1 + pi/2)r You forgot to divide 31 by 2. otherwise your work is fine.

OpenStudy (phi):

A box is to be made out of a 8 by 16 piece of cardboard. Squares of equal size will be cut out of each corner, and then the ends and sides will be folded up to form a box with an open top. Find the length L, width W, and height H of the resulting box that maximizes the volume. (Assume that W≤L).  replace W and L in the equation for the volume and take the derivative with respect to H